# FMCW Radar Principles and Human Activity Recognition Systems: Foundations, Techniques, and Applications

### v1

@@ 27.1 Introduction to Transformer Models
Transformer models were initially developed for natural language processing (NLP) tasks, such as language translation and text generation. However, their highly flexible architecture, specifically the self-attention mechanism [64], makes them well-suited for various other tasks, including radar-based Human Activity Recognition (HAR). Transformers excel at handling temporal sequences and capturing both local and global dependencies within the data, which are crucial for accurately identifying and classifying human activities based on radar signals.
@@ 27.2.1 How Self-attention Works
In radar-based HAR, this means that the Transformer can compare different time frames of radar data to determine which parts of the sequence are most relevant for identifying specific activities. This capability allows the model to capture long-term temporal dependencies [24], even if the activity consists of subtle motions over a longer period.
@@ 27.2.2 Parallel Processing and Temporal Flexibility
One key advantage of the Transformer’s self-attention mechanism is its ability to process data in parallel [26]. Instead of sequentially processing each time frame like LSTM models, Transformers can process the entire sequence at once, allowing for faster training and inference.
@@ 27.3.1 Example: Radar Data with Transformer
4class RadarTransformer(nn.Module):
@@
7 self.encoder_layer = nn.TransformerEncoderLayer(d_model=num_features, nhead=num_heads)
8 self.transformer = nn.TransformerEncoder(self.encoder_layer, num_layers=num_layers)
9 self.fc = nn.Linear(num_features, 10) # Assume 10 classes for activity recognition
@@ 27.3.7 PointNet for Point Cloud Data
For radar systems that generate point cloud data (e.g., 3D radar), PointNet can be used to extract spatial features from the point clouds [41]. PointNet is a neural network architecture designed specifically to process point clouds and is well-suited for radar systems that track multiple targets in a 3D space.
@@
1class CNNTransformer(nn.Module):
@@
5 self.cnn = nn.Conv2d(1, 32, kernel_size=3, stride=1, padding=1)
6 self.pool = nn.MaxPool2d(2, 2)
7 self.encoder_layer = nn.TransformerEncoderLayer(d_model=num_features, nhead=num_heads)
8 self.transformer = nn.TransformerEncoder(self.encoder_layer, num_layers=num_layers)
9 self.fc = nn.Linear(num_features, 10) # Assume 10 classes for activity recognition
@@ 29.2 Overview of Signal Processing Techniques
Signal processing plays a vital role in extracting meaningful information from raw radar signals [44]. Several key steps transform the radar data into interpretable results, laying the foundation for tasks like target detection and human activity recognition.
@@ 29.3 Deep Learning for Radar Data
Transfer Learning and Pre-trained Models: Transfer learning enables the use of pre-trained models [56] (trained on large datasets such as ImageNet or ModelNet) to be fine-tuned on radar-specific tasks. This helps in reducing training time and improving performance, especially when the available radar data is limited.
@@
9model.fc = nn.Linear(model.fc.in_features, 10)
@@ 29.4 Future Directions in Radar-based HAR Systems
Looking ahead, several exciting directions are emerging in the field of radar-based human activity recognition [33] (HAR):

### v2

@@ 27.1 Introduction to Transformer Models
Transformer models were initially developed for natural language processing (NLP) tasks, such as language translation and text generation. However, their highly flexible architecture, specifically the self-attention mechanism [wei2021self], makes them well-suited for various other tasks, including radar-based Human Activity Recognition (HAR). Transformers excel at handling temporal sequences and capturing both local and global dependencies within the data, which are crucial for accurately identifying and classifying human activities based on radar signals.
@@ 27.2.1 How Self-attention Works
In radar-based HAR, this means that the Transformer can compare different time frames of radar data to determine which parts of the sequence are most relevant for identifying specific activities. This capability allows the model to capture long-term temporal dependencies [huang2023long], even if the activity consists of subtle motions over a longer period.
@@ 27.2.2 Parallel Processing and Temporal Flexibility
One key advantage of the Transformer’s self-attention mechanism is its ability to process data in parallel [jain2021super]. Instead of sequentially processing each time frame like LSTM models, Transformers can process the entire sequence at once, allowing for faster training and inference.
@@ 27.3.1 Example: Radar Data with Transformer
4class RadarTransformer($n$.Module):
@@
7 self.encoder_layer = $n$.TransformerEncoderLayer(d_model=num_features, nhead=num_heads)
8 self.transformer = $n$.TransformerEncoder(self.encoder_layer, num_layers=num_layers)
9 self.fc = $n$.Linear(num_features, 10) # Assume 10 classes for activity recognition
@@ 27.3.7 PointNet for Point Cloud Data
For radar systems that generate point cloud data (e.g., 3D radar), PointNet can be used to extract spatial features from the point clouds [paigwar2019attentional]. PointNet is a neural network architecture designed specifically to process point clouds and is well-suited for radar systems that track multiple targets in a 3D space.
@@
1class CNNTransformer($n$.Module):
@@
5 self.cnn = $n$.Conv2d(1, 32, kernel_size=3, stride=1, padding=1)
6 self.pool = $n$.MaxPool2d(2, 2)
7 self.encoder_layer = $n$.TransformerEncoderLayer(d_model=num_features, nhead=num_heads)
8 self.transformer = $n$.TransformerEncoder(self.encoder_layer, num_layers=num_layers)
9 self.fc = $n$.Linear(num_features, 10) # Assume 10 classes for activity recognition
@@ 29.2 Overview of Signal Processing Techniques
Signal processing plays a vital role in extracting meaningful information from raw radar signals [pham2021intelligent]. Several key steps transform the radar data into interpretable results, laying the foundation for tasks like target detection and human activity recognition.
@@ 29.3 Deep Learning for Radar Data
Transfer Learning and Pre-trained Models: Transfer learning enables the use of pre-trained models [sohail2024advancing] (trained on large datasets such as ImageNet or ModelNet) to be fine-tuned on radar-specific tasks. This helps in reducing training time and improving performance, especially when the available radar data is limited.
@@
9model.fc = $n$.Linear(model.fc.in_features, 10)
@@ 29.4 Future Directions in Radar-based HAR Systems
Looking ahead, several exciting directions are emerging in the field of radar-based human activity recognition [li2019survey] (HAR):